OpenStudy (anonymous):

Can someone help me with: z^3 = 27i ?

OpenStudy (anonymous):

What is the symbol after Z supposed to be?

OpenStudy (noorik.):

you need to find z value, don't you?

OpenStudy (anonymous):

elevated in third Pepitop

OpenStudy (anonymous):

Yes NooriK

OpenStudy (noorik.):

do you know De Moivre's theorem?

OpenStudy (anonymous):

yes I know, but I can't figure out what to do..

OpenStudy (noorik.):

first you have to find mod (z) mod(z)=sqrt(x^2+y^2) mod(z)=sqrt(0+27^2) then \[\cos \theta = \frac{ x }{ \mod(z) }\] \[\sin \theta = \frac{ x }{ \mod(z) }\] then \[\arg(z) = \theta\] then \[z^3 = mod(z) * (\cos\theta +\sin\theta)\] then \[z =\sqrt[3]{\mod(z)}*(\cos \frac{ \theta+2k \pi }{ 3 }+i \sin \frac{ \theta+2k \pi }{ 3 })\] it's hard, i know that, but i hope that'll help

OpenStudy (anonymous):

Thanks! Very helpful

OpenStudy (noorik.):

you must note that, in the last form its cos ((mod(z)+2piK)/3)+i sin((mod(z)+2piK)/3)

OpenStudy (noorik.):

you are welcome

OpenStudy (anonymous):

What is θ here?

OpenStudy (noorik.):

you know that we have transformed it into the polar form, so theta is an angle

OpenStudy (anonymous):

yeah I know but how do you find it, i'm very newbie at this kind of math..